# FedRef: Communication-Efficient Bayesian Fine Tuning with Reference Model

### v1

@@ Abstract
Federated learning(FL) is used for distributed scenarios to train artificial intelligence(AI) models while ensuring users’ privacy. In federated learning scenario, the server generally never knows about users’ data. This type of concept makes the AI training process efficient in terms of data privacy. However, regarding model performance, federated AI models may not sufficiently satisfy AI users’ expectations. Furthermore, AI users have a wide range of different needs. It is not easy to satisfy the whole users needs. These types of issues can be addressed through AI model optimization, fine-tuning, or personalization to achieve optimal model performance. To address model optimization challenges, we propose reference model-based federated learning for optimal fine-tuning, which overcomes catastrophic forgetting in each round. This method is derived from Bayesian parameter-efficient transfer learning, which includes an optimal proximal term and enables overcoming the catastrophic forgetting issue in each round by utilizing a reference model that incorporates previous model parameters. As a result, this method achieves both high model performance and low computing cost.
@@ I Introduction
In the related work section, we introduce existing approaches (FedAvg, FedProx, FedOpt, Bayesian Transfer Learning) for model optimization in FL scenarios, along with their limitations regarding client computing cost and catastrophic forgetting [2][3] [4] in each round. We then introduce our FedRef: a communication-efficient Bayesian fine-tuning approach with a reference model. This concept overcomes catastrophic forgetting in each round by integrating previous model features into the maximum a posteriori (MAP) problem.
@@
In our experiments, we applied FedRef on MNIST, CIFAR10, FeTs2022 data partitioning from distributed condition for non-IID(Independent and Identically Distributed) setting. We defined each baseline and then analyzed model performances, as baseline models differ for each dataset. Especially, on the FeTs2022, the data is derived in medical imaging area for brain tumor segmentation. In this case, the client-side can be a hospital, as shown in Fig. 1.
Figure 1: Basic federated learning system.
As a result, our FedRef achieves high model performance, despite minimizing clients’ computing cost.
@@ II-1 FedAvg
$s$ : number of local steps for local training, $\eta$: learning rate, $\mathcal{B}_{i}^{(s)}$: mini-batch sampled from $i$ index client local dataset on step $s$, $F_{i}(\theta_{i}^{(s)};\mathcal{B}_{i}^{(s)})$: local loss function evaluated on a mini-batch. The local training in our FedRef client update is based on the SGD solution. In algorithm 1, it show whole FedAvg process flow.
@@ II-3 FedOpt
$v_{t}=v_{r-1}-(1-\beta_{2})\cdot sign(v_{r-1}-g_{r}^{2})\cdot g_{r}^{2},$
@@ III FedRef: Communication-Efficient Bayesian Fine Tuning with Reference Model
In equation (III), the constant value $\sum_{k}^{K}F_{k}$ denotes the sum of client losses. The diagonal matrix $\mathrm{diag}(W_{1},\dots,W_{K})$ represents aggregation weights (e.g. $\frac{n_{i}}{n}$). $\sum_{i}(\theta_{i}-\theta_{0,i})^{2}$ signifies $L_{2}$ regularization of model $\theta-\theta_{0}$. Regarding parameter requirements, only clients’ losses are needed for our FedRef concept. Only on the server side, model optimization is performed, which can decrease both client computing cost and communication cost. In the FedRef, $\theta^{2}$ should be set as the reference model $\theta_{ref}$. For detail on $\theta_{ref}$, we defined the reference model as shown in equation 13.
@@
Experimentally, a suitable value for $p$ was analyzed to be 5. For the whole algorithm, we denoted FedRef at following algorithm5:
@@ IV-A1 Global
Within the Python language environment, the Flower framework enables easy construction of federated learning servers and clients from high-level code, based on Object-Oriented Programming (OOP) and Functional Programming (FP), for privacy-preserving machine learning strategies(e.g. FedAvg, FedProx, FedOpt, etc.).
@@ IV-A2 MNIST
IV-A3 CIFAR10
The CIFAR-10 (Canadian Institute for Advanced Research) dataset[18] is one of the most prominent and widely used image datasets in machine learning and computer vision research. It consists of 60,000 32x32 color images across 10 distinct classes, making it a popular benchmark for evaluating the performance of classification algorithms, particularly in the realm of image recognition. The data partitioning mechanism for this dataset is the same as the MNIST setting. For the model baseline, ResNet [19] is applied for CIFAR10 classification task.
IV-A4 FeTs2022
The FeTs2022 dataset refers to the dataset used in the Federated Tumor Segmentation (FeTS) Challenge 2022, which focused on brain tumor segmentation using federated learning approaches. The data includes multimodal brain MRI scans of glioma patients. Each patient folder typically has the four MRI modalities (T1-weighted, T1-weighted post-contrast, T2-weighted, T2 Fluid Attenuated Inversion Recovery). The data labels include segmentations of three tumor sub-regions (Enhancing tumor, Tumor core, Whole tumor). The data can be downloaded at official Synapse site [20]. We perform the segmentation task for evaluating FedRef on FeTs2022 data which have specific non-IID data partitioning at Fig. 2.
@@ IV-B Result
Figure 3: Graph of centralized loss on each strategy and data
TABLE II: Performance comparison on segmentation task from FeTs2022
@@
FedOpt
FedRef

### v2

@@ Abstract
Federated learning(FL) is used for distributed scenarios to train artificial intelligence(AI) models while ensuring users’ privacy. In federated learning scenario, the server generally never knows about users’ data. This type of concept makes the AI training process efficient in terms of data privacy. However, regarding model performance, federated AI models may not sufficiently satisfy AI users’ expectations. Furthermore, AI users have a wide range of different needs. It is not easy to satisfy the whole users needs. These types of issues can be addressed through AI model optimization, fine-tuning, or personalization to achieve optimal model performance. To address model optimization challenges, we propose reference model-based federated learning for optimal fine-tuning, which overcomes catastrophic forgetting in each round. This method is derived from Bayesian parameter-efficient transfer learning, which includes an optimal proximal term and utilizes a reference model that incorporates previous model parameters. As a result, this method achieves both high model performance and clients’ low computing cost.
@@ I Introduction
In the related work section, we introduce existing approaches (FedAvg, FedProx, FedOpt, Bayesian Transfer Learning) for model optimization in FL scenarios, along with their limitations regarding client computing cost and catastrophic forgetting [2][3] [4] issue. We then introduce our FedRef: a communication-efficient Bayesian fine-tuning approach with a reference model. This concept overcomes catastrophic forgetting in each round by integrating previous model features into the maximum a posteriori (MAP) problem.
@@
In our experiments, we applied FedRef on MNIST, DRIVE, FeTs2022, BRATS data partitioning from distributed condition for non-IID(Independent and Identically Distributed) setting. We defined each baseline and then analyzed model performances, as baseline models differ for each dataset. Especially, on the FeTs2022 and BRATS, the dataset are derived in medical imaging area for brain tumor segmentation. In this case, the client-side can be a hospital, as shown in Fig. 1. As a result, our FedRef achieves high model performance, despite minimizing clients’ computing cost in medical area and general computer vision tasks. In the Result section, we represented models performance of each round by graph and specifically, the FeTs2022 task table shows specific evaluation metrics (mean Intersection of Union, Dice Coefficient, Hausdorff Distance) [5][6][7][8] which generally be used on medical segmentation tasks.
Figure 1: Basic medical federated learning system.
@@ II-1 FedAvg
$s$ : number of local steps for local training, $\eta$: learning rate, $\mathcal{B}_{i}^{(s)}$: mini-batch sampled from $i$ index client local dataset on step $s$, $F_{i}(\theta_{i}^{(s)};\mathcal{B}_{i}^{(s)})$: local loss function evaluated on a mini-batch. In algorithm 1, it show whole FedAvg process flow.
@@ II-3 FedOpt
$v_{r}=v_{r-1}-(1-\beta_{2})\cdot sign(v_{r-1}-g_{r}^{2})\cdot g_{r}^{2},$
@@ III FedRef: Communication-Efficient Bayesian Fine Tuning with Reference Model
In equation (III), the constant value $\sum_{k}^{K}F_{k}$ denotes the sum of client losses. The diagonal matrix $\mathrm{diag}(W_{1},\dots,W_{K})$ represents aggregation weights (e.g. $\frac{n_{i}}{n}$). $\sum_{i}(\theta_{i}-\theta_{0,i})^{2}$ signifies $L_{2}$ regularization of model $\theta-\theta_{0}$. Regarding parameter requirements, only clients’ losses are needed for our FedRef concept. Only on the server side, model optimization is performed, which can decrease client computing cost. In the FedRef, $\theta^{2}$ should be set as the reference model $\theta_{ref}$. For detail on $\theta_{ref}$, we defined the reference model as shown in equation 13.
@@
Experimentally, a suitable value for $p$ was analyzed heuristically to be 3 to 5, but it should be considered about memory resources for saving parameters. For the whole algorithm, we denoted FedRef at following algorithm5:
@@ IV-A1 Global
Within the Python language environment, the Flower framework enables easy construction of federated learning servers and clients from high-level code, based on Object-Oriented Programming (OOP) and Functional Programming (FP), for privacy-preserving machine learning strategies(e.g. FedAvg, FedProx, FedOpt, etc.). The source code for our method is publicly available at [https://github.com/TaehwanY98/Fed-Ref]
@@ IV-A2 MNIST
IV-A3 FeTs2022
The FeTs2022 dataset refers to the dataset used in the Federated Tumor Segmentation (FeTS) Challenge 2022, which focused on brain tumor segmentation using federated learning approaches. The data includes multimodal brain MRI scans of glioma patients. Each patient folder typically has the four MRI modalities (T1-weighted, T1-weighted post-contrast, T2-weighted, T2 Fluid Attenuated Inversion Recovery). The data labels include segmentations of three tumor sub-regions (Enhancing tumor, Tumor core, Whole tumor). The data can be downloaded at official Synapse site [22]. We perform the segmentation task for evaluating FedRef on FeTs2022 data which have specific non-IID data partitioning at Fig. 2. Baseline model is used to 3D U-Net.
IV-A4 Drive
The DRIVE dataset is a widely recognized resource in the field of retinal image analysis, particularly for blood vessel segmentation. It helps researchers study, diagnose, and manage various cardiovascular and ophthalmic diseases like diabetes and hypertension by analyzing morphological properties of retinal vessels such as length, width, tortuosity, and branching patterns. The dataset comprises 40 color fundus photographs. Seven of these images show signs of mild early diabetic retinopathy, representing abnormal pathological cases. All images were collected from a diabetic retinopathy screening program in the Netherlands [23].
@@ IV-B Result
Figure 3: Graph of server-side centralized loss on each strategy and data.
Figure 4: Graph of server-side centralized Dice-coefficient(DCE) and Hausdorff distance. Left: DCE, Right: HF95
TABLE II: Performance comparison from Medical data
@@
FedOpt 16.67% 21.39% 12.54
FedRef 16.81% 21.53% 12.13
Fed-Avg 30.23% 45.95% 14.82
[…]
FedOpt 30.23% 45.95% 14.82
FedRef 32.27% 48.17% 14.79